OpenStudy (anonymous):

Factorise f(x)= x^6+x^4+x^2+1 by grouping in pairs. Hence find all the roots of f(x) = 0, using complex numbers.Please do not use pie as i havent learnt that yet :)

OpenStudy (anonymous):

Write it as a geometric series under the substitution u:x^2, then factor it into roots of unity.

OpenStudy (psymon):

Lol, wtf? Bit much for this.

OpenStudy (anonymous):

how is that a bit much, I don't see any other way of doing it

OpenStudy (anonymous):

i get how to factorise by grouping in pairs i get (x^2+1)(x^4+1), then what do i do?

OpenStudy (anonymous):

Factor each of those expressions

OpenStudy (psymon):

Set each factor equal to 0. x^2 + 1 = 0 x^4 + 1 = 0

OpenStudy (anonymous):

There going to be roots of unity

OpenStudy (anonymous):

Of the form [\e^{2\pi i t}\]

OpenStudy (anonymous):

The first one has roots $$-i,i$$

OpenStudy (anonymous):

yeah for x^2+1=0, i got x= -i and x=i, and i havent learn that stuff with pie yett and for x^4 + 1 = 0 i get square root of i, and sqaure root of -i

OpenStudy (anonymous):

You try the second one

OpenStudy (psymon):

That's all ya gotta do. Your zeros are +/- i.

OpenStudy (anonymous):

but the answer says \[\pm(1/\sqrt{2} + 1/\sqrt{2} i ) , \pm i, \pm(1/\sqrt{2} - 1/\sqrt{2} i )\]

OpenStudy (psymon):

Hmm.....maybe this is something I haven't learned yet. Is just regular algebra or higher level?

OpenStudy (psymon):

@rebeccapecovski I looked it up, guess it isn't something I would have thought of. I have no idea about the pi thing either you mentioned, but I do have a way to go about showing the answer that has nothing to do with pi.

OpenStudy (luigi0210):

I think I seeyour mistake, you factored wrong

OpenStudy (psymon):

Nah, she factored correctly, it's just a bit.....more complicated than that, lol.

OpenStudy (luigi0210):

You sure about that? I may be wrong but just making sure.. it's 1 AM right now so

OpenStudy (psymon):

Well, what do you think then?

OpenStudy (luigi0210):

I can't think at the moment, my brain is on an extended vacation -_-

OpenStudy (psymon):

Haha. Well, I would say that you have \[\pm i, \pm \sqrt{i}\] and then getting +/- sqrt i is it's own adventure.

hero (hero):

x^6 + x^4 + x^2 + 1 = 0 x^4(x^2 + 1) + 1(x^2 + 1) = 0 (x^2 + 1)(x^4 + 1) = 0 x^2 + 1 = 0 x^4 + 1 = 0 x^2 = -1 x^4 = -1 x = +/- sqrt(-1) x + +- qdrt(-1)